OpenStudy (anonymous):

Solve w + 12 = –6 – 2w.

OpenStudy (eyust707):

Steps: Bring all the terms with w to the same side. Bring all the terms with numbers to the other side. Divide both sides by coefficient in front of the w.

OpenStudy (anonymous):

get the common numbers together:12+6=-w-2w 18=2w^2 9=w... i think thats right :))))